OpenStudy (anonymous):

Julien & Stacy needed 10 liters of water to fill a tank. Stacy filled the tank with 3 11/12 liters of water. Julien poured 1 2/5 liters less than Stacy in the tank. How much more water is still needed to fill the tank? HELP!

OpenStudy (xkehaulanix):

If you know how many liters Stacy poured into the tank, you can find how much Julien poured in. Stacy poured in 3 and 11/12 liters, and Julien poured 1 and 2/5 less, so Julien poured \[3+11/12 -(1 + 2/5)\] To add these, you have to find common denominators \[(36/12 + 11/12) - (5/5 + 2/5)\]\[=47/12-7/5\]\[=235/70 - 84/5\]\[=151/5\] So Julien poured that much. Add it to what Stacy poured into the tank to get the total amount of water poured, then find the difference between how much was poured and how much you need.